OpenStudy (anonymous):

x=-b+/- multipled by the square root of b squared - 4 ac

OpenStudy (anonymous):

yes?

myininaya (myininaya):

\[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

ok look my teacher gave us some equations and one of them are \[{x}^2+6x+9=0\] and the answeres are x=-3 so i dont get it

OpenStudy (anonymous):

yes, this is (x+3)^2 so x = -3

OpenStudy (anonymous):

ok ;) so how do iget the answer for the first one

OpenStudy (anonymous):

Take the coefficient of the terms: i.e. the function is in ax^2+bx+c=0 form and plug the values of a, b, and c into the formula myiniaya posted, that is one way.

OpenStudy (anonymous):

which one is the first one?

OpenStudy (anonymous):

the first one is \[x=-b \pm \sqrt{b ^{2}- 4ac} \div 2a\] ;D

myininaya (myininaya):

what do you want to solve for?

OpenStudy (anonymous):

what about it?

myininaya (myininaya):

we can solve it for b, a, or c? its already solved for x

OpenStudy (anonymous):

that is the quadratic formula to find roots.

OpenStudy (anonymous):

well yeah the teacher wants us to find all thee a b and c but i dont reallly know how!

myininaya (myininaya):

she wants you to solve the quadratic formula for a then b then c are you sure?

OpenStudy (anonymous):

yes thats what she said :S

myininaya (myininaya):

k let's see if we can do it let's try to solve for a first

OpenStudy (anonymous):

ok ;D

OpenStudy (anonymous):

it is impossible to solve this equation for a b and c

myininaya (myininaya):

one sec

myininaya (myininaya):

i think you can use the quadratic formula to solve for a lol

OpenStudy (anonymous):

which is??

OpenStudy (anonymous):

yes, but it impossible to solve for all three values of a,b and c

myininaya (myininaya):

lol. i honestly don't think this what your teacher wants

myininaya (myininaya):

i can solve for b and c use the quad on b and c is easiest

OpenStudy (anonymous):

no, you cannot. x can hold only two values. you need atleast 3 equations to solve for 3 unknown variables.

myininaya (myininaya):

a doeesnt have to be both it can be one or the other just like x

OpenStudy (anonymous):

you may be able to get a,b and c in terms of the other variables, but that doesn't get you any closer to solving for them.

myininaya (myininaya):

right but it sounded like he wanted us to solve the quadratic formula for each lol

OpenStudy (anonymous):

oh, that is just an exercise in stupidity.

OpenStudy (anonymous):

there is no point of that. I don't think their teacher asked them to solve for a,b and c.

myininaya (myininaya):

me neither